OpenStudy (jensine):

sovle the equation 180a^2+300a+125=0

OpenStudy (anonymous):

solve for a?

OpenStudy (agreene):

180a^2+300a+125=0 (a+5/6)^2=0 Factor Left Side a+5/6=0 sqrt both sides a=-5/6 -5/6 both sides

OpenStudy (jensine):

how do u do this problem? 2x^2+28x+98=0

OpenStudy (anonymous):

great work @Luis_Rivera